# Credit Risk Meets Large Language Models: Building a Risk Indicator from Loan Descriptions in P2P Lending

### v2

@@ 5.2.2 Avoiding data leakage in cross-validation
Figure 1: BERT training diagram.
@@ 5.2.3 Result of the transfer learning process
5.2.4 Inclusion of the BERT score in the classification model
The experiment architecture is shown in Figure 2. First, the loan descriptions are processed by the fine-tuned BERT model. By using a sigmoid activation function in its final layer, the resulting output can be interpreted as the probability of belonging to the positive class. This probability, denoted as BERT_score, represents a default score solely based on the content of the loan description. Finally, this BERT score is included as an input variable in the XGBoost classifier along with quantitative and categorical variables to generate a final default prediction.
Figure 2: Diagram of the experiment architecture.
@@ 6.1 Assessment of the BERT score as a credit risk score
6.2 Assessment of the relationship between the BERT score and other variables
In this section, we aim to explore potential relationships between the BERT score and other variables. Table 9 presents the correlation coefficients between the BERT score and the quantitative variables within the dataset. The findings reveal weak yet statistically significant relationships with all variables except for the loan amount. Notably, the most pronounced correlations exist with the FICO score and revenue variables. Both exhibit inverse relationships, indicating that individuals with higher FICO scores and revenues tend to have lower BERT scores, and vice versa. The association between the BERT score and the debt variable (dti_n) is direct, albeit slightly weaker than the other two correlations.
@@ 6.2 Assessment of the relationship between the BERT score and other variables
To evaluate the relationship with the categorical variables, Table 10 presents the results of the Kruskal-Wallis test. The Kruskal-Wallis test is a non-parametric statistical test that analyzes whether there are statistical differences in the BERT scores across categories within each categorical variable. The results indicate significant differences in BERT scores among all categorical variables, suggesting a certain level of association between the BERT score and these categorical factors. However, it remains challenging to quantify the strength of this relationship or identify the specific categories with the most robust associations.
@@ 7.1 Analysis of the classification performance
First, we evaluate the impact of incorporating the BERT score in the granting model. In our baseline experiment, we optimize XGBoost with a genetic algorithm using the quantitative and categorical variables typically used in granting models, while in the competing approach, we optimize it but include the BERT score as input variable. Table 12 shows the results of both approaches.
A closer examination of the balanced metrics reveals a marginal enhancement in both BACC (0.6154 vs. 0.6187) and AUC (0.6575 vs. 0.6644), the latter significant according to the DeLong test [46] at the 0.01 level. It is crucial to note that the Lending Club dataset exclusively consists of approved loans. This fact poses a substantial challenge to significantly enhance the outcomes in a loan granting model such as ours since the loans included in the dataset were initially considered favorable by the platform. Furthermore, in experiments not reported in the paper we used CatBoost [47] instead of XGBoost and obtained a similar BACC improvement.
@@
Table 13 shows an additional experiment in which XGBoost classifiers are trained and optimized using only one kind of input variable. While the classifier using the quantitative variables is clearly the best, the classifier that uses just the BERT score obtains slightly better results than the one using the four categorical variables (the AUC difference is significant at the 0.01 level according to the DeLong test). This result is noteworthy given the well-known effectiveness and the meaningful nature of the qualitative variables.
@@ 7.2 Feature importance and explainability
To further analyze the impact of the BERT score, in this section, we analyze whether the models’ use of the input variables changes. Figure 4 shows the feature importance given by XGBoost both with and without incorporating the BERT score. Notably, the BERT score emerges as the third most influential variable, accounting for 6% importance. Moreover, the order and magnitude of importance for other variables exhibit shifts. Specifically, the borrower’s annual revenue diminishes in importance following the inclusion of the BERT score, while the significance of the ‘credit card’ purpose increases. This observation underscores the significant influence of the BERT score on the classifier.
@@
Figure 4: Feature importances of the XGBoost classifiers.
To gain a deeper understanding of the role of variables, we employ SHAP values [48]. Figure 5 provides a visualization of the SHAP values for the 10 variables with the most significant impact on the model’s output, comparing models with and without including the BERT score.
@@
It is noteworthy that Figure 5(a) resembles a comparable figure presented in [41], wherein SHAP values were utilized in a lending model on Lending Club [^11] Our dataset is more constrained, only considering loans with textual descriptions, as explained in Section 4.. Examining Figure 5(b), we find that the BERT score ranks as the fourth variable in terms of impact. A closer inspection of the violin plots reveals that including BERT does not substantially alter how variables influence the target class.
@@
Figure 6 reveals a direct and linear relationship between the BERT score and the default risk. Notably, this relationship is asymmetric around BERT score values of 0.5; scores below 0.4 correspond to SHAP values ranging between -0.4 and -0.6, leading the model to predict non-default. Conversely, this impact range in the positive case is only reached by BERT scores exceeding 0.7, signifying that only exceptionally high BERT scores serve as strong indicators of default.
@@
7.3 Impact across the purpose categories
@@
Now we delve into the influence of incorporating the BERT score on classification outcomes across various categories of the purpose variable. Leveraging the language comprehension capabilities of LLMs, it is conceivable that the BERT model can provide more accurate characterizations of loan purposes than the purpose variable alone. For instance, the ‘other’ category, being inherently ambiguous, could benefit from a more nuanced understanding of the loan description, potentially leading to improved prediction outcomes.
@@
Figure 7: BACC changes by purpose after including the BERT score.
@@
Figure 7 illustrates the changes in balanced accuracy upon incorporating the BERT score for each loan purpose. Notably, the ‘other’ category shows a modest improvement of 2.11%. However, several other categories exhibit more substantial enhancements, including ‘educational’ (9.22%), ‘moving’ (5.84%), ‘medical’ (3.84%), and ‘small business’ (3.40%). Given the black-box nature of our model, it is difficult to ascertain the reasons why these categories have improved more than the others. However, we posit that these categories share a commonality—the more detailed specification of purpose or a deeper understanding of the borrower’s situation contributes to a more precise delineation of the default risk. For instance, in education loans, factors like the field of study and the educational institution may play a crucial role in determining employability and salaries, thereby enabling a more accurate assessment of the default risk. The same happens in the case of moving to a new place, having to pay medical expenses, or obtaining funds for a business.
@@
The following examples illustrate instances of loans that were initially misclassified as defaults in the absence of the BERT score but were subsequently correctly predicted as non-defaults after its inclusion:
@@
In each of these examples, the corresponding BERT scores fall below 0.4. As demonstrated in Figure 6, the classifier will have a stronger driver to categorize the loan as non-default when the score is below this value. As all the examples offer more details on the situation and the purpose of the loan, their further clarity could explain their low BERT score. In addition, all the texts shown are accurate, readable, and written in a confident and objective tone. While previous evidence in the literature offered mixed results about the influence of linguistic factors as default predictors [3, 18, 4, 7], we have shown that our BERT score is significantly related to them. While these explanations are sound, given the black-box nature of our BERT-based approach, the explanations influencing the BERT score of each text cannot be precisely determined. This aspect needs to be improved to comply with the requirements of the regulators and foster trust among the end-users.
8 Conclusions
In this paper, we have introduced a novel approach that highlights the value of state-of-the-art natural language processing techniques in enhancing credit risk models. Our approach fine-tunes BERT using loan descriptions to generate a risk score that effectively discriminates between defaulted and non-defaulted loans. Furthermore, we show that integrating the BERT-based risk score with traditional variables significantly enhances the performance of conventional granting models. These results underscore the potential of LLMs to augment lending models, emphasizing their invaluable role in harnessing the richness of textual data to improve credit risk assessment.
Our approach can be easily applied without manual annotation, which is a complex task that also introduces annotator subjectivity. As our analysis shows, the score incorporates both linguistic factors and content aspects that relate to the creditworthiness of the borrower’s situation and purpose. Additionally, while fine-tuning the LLM is a time-consuming process that requires GPU processing, generating predictions, such as the risk score for a loan description, can be done instantly with standard home equipment.
Our work represents just the first steps in this endeavor and underscores the necessity for further exploration and integration of cutting-edge methodologies to enhance the effectiveness of risk-return assessments in the dynamic landscape of P2P lending or even in traditional financial institutions. Several directions deserve to be explored to refine the predictive capability of the model and the understanding of loan applicants’ conditions.
An important limitation of our work is the lack of transparency regarding the factors contributing to a given risk score and its potential biases. These issues hinder its application in real-world contexts, where a comprehensive understanding of the score generation process and assurance that it does not introduce biases related to gender, ethnicity, financial inclusion, or other factors are essential. Thus, efforts to enhance the transparency and explainability of the risk score are crucial not only for regulatory compliance but also for fostering trust among borrowers and lenders [49].
@@ 8 Conclusions
At this respect, LLM-based topic modeling tools such as BERTopic [50] could be used to gain an understanding of the risk scores. For example, topic modeling could identify whether a loan description is related to “risky topics” that in turn could be incorporated as input variables in the granting model. Similarly, the embeddings from BERT or other LLMs could be used to generate interpretable topics from large, complex vocabularies, as other works have done [51]. Applying such topic modeling approaches to loan descriptions could facilitate the identification of situations with varying risk levels. These topics could then serve as additional variables in classification models, improving predictive performance while maintaining the transparency required in financial assessments.
@@
Another potential research avenue involves delving into more advanced LLMs. For instance, encoder-only models like RoBERTa [37] could improve the model’s capacity to capture more intricate linguistic patterns. Furthermore, the rise of generative AI has spurred interest in exploring decoder models for classification purposes. Although these models were not originally designed for such tasks, recent advancements have shown promising approximations. For example, CARP (Clue And Reasoning Prompting) [52] utilizes in-context learning, generating textual responses conditioned on a provided prompt with a few annotated examples (few-shot learning). Remarkably, this approach improved results in classification tasks without the need for transfer learning and fine-tuning and could be applied to risk scoring.
@@
Future work should also address the economic implications of our findings by integrating cost- or profit-sensitive approaches [53, 42]. These methods allow for the inclusion of the financial impacts of classification outcomes. Investigating whether improvements in model performance by including textual descriptions translate into better financial outcomes would enhance the real-world applicability of our findings.

### v3

@@ 5.2.2 Avoiding data leakage in cross-validation
Figure 3: Data handling strategy to prevent data leakage in model training and evaluation.
@@ 6.1 Assessment of the BERT score as a credit risk score
6.2 Relationship between the BERT score and other variables
We now explore potential relationships between the BERT score and other variables. Table 9 presents the correlation coefficients between the BERT score and the quantitative variables within the dataset. The findings reveal weak yet statistically significant relationships with all variables except for the loan amount. Notably, the most pronounced correlations exist with the FICO score and revenue variables. Both exhibit inverse relationships, indicating that individuals with higher FICO scores and revenues tend to have lower BERT scores, and vice versa. The association between the BERT score and the debt variable (dti_n) is direct, albeit slightly weaker than the other two correlations.
@@ 6.2 Relationship between the BERT score and other variables
To evaluate the relationship with the categorical variables, Table 10 presents the results of the Kruskal-Wallis test, a non-parametric statistical test that analyzes whether there are differences in the BERT scores across categories within each categorical variable. The results indicate significant differences in BERT scores among all categorical variables, suggesting a certain level of association between the BERT score and these categorical factors. However, it remains challenging to quantify the strength of this relationship or identify the specific categories with the most robust associations.
@@ 7.1 Analysis of the classification performance
First, we evaluate the impact of incorporating the BERT score in the granting model. In our baseline experiment, we optimize XGBoost with a genetic algorithm using the quantitative and categorical variables typically used in granting models, while in the competing approach, we optimize it but include the BERT score as an input variable. Table 12 shows the results of both approaches.
A closer examination of the balanced metrics reveals a marginal enhancement in both BACC (0.6154 vs. 0.6187) and AUC (0.6575 vs. 0.6644), the latter significant according to the DeLong test [50]—a non-parametric approach for evaluating whether differences between AUCs of two models are statistically significant— at the 0.01 level. It is crucial to note that the Lending Club dataset exclusively consists of approved loans. This fact poses a substantial challenge to significantly enhance the outcomes in a loan granting model such as ours since the loans included in the dataset were initially considered favorable by the platform. Furthermore, in experiments not reported in the paper we used CatBoost [37] instead of XGBoost and obtained a similar BACC improvement.
@@
Table 13 shows an additional experiment in which XGBoost classifiers are trained and optimized using only one kind of input variable. While the classifier using the quantitative variables is clearly the best, the classifier that uses just the BERT score obtains slightly better results than the one using the four categorical variables (the AUC difference is significant at 0.01 according to the DeLong test). This is noteworthy given the well-known effectiveness and the meaningful nature of the qualitative variables.
@@ 7.2 Feature importance and explainability
Now we delve into how the inclusion of the BERT score alters the classifier’s use of input variables. Our goal is not only to assess the predictive contribution of the BERT score but also to understand its broader impact on the model’s decision-making process. Figure 5 shows the feature importance assigned by XGBoost, both with and without the BERT score. Notably, the BERT score becomes the third most influential variable, accounting for 6% of the total importance. Moreover, its inclusion reshapes the importance and ranking of other features. For instance, the importance of the borrower’s annual revenue and the debt-to-income ratio (dti_n), both crucial for assessing the borrower’s economic status, significantly decreases. This shift suggests that the BERT score may encapsulate information that overlaps with these variables or renders them less critical, as supported by the correlation analysis in Table 9.
@@
Figure 5: Feature importance of the two XGBoost classifiers.
@@
To analyze whether the relationships between each variable and the predicted outcome change, we use the SHAP values [30], which quantify the contribution of each feature to the model’s predictions. Figure 6 compares the SHAP values for the 10 most impactful features in models both with and without the BERT score. It reveals that the distributions of SHAP values, as seen in the violin plots, do not significantly change when the BERT score is included. Interestingly, the SHAP value distribution in Figure 6(a) mirrors a similar analysis presented in [2], conducted on a lending model with Lending Club data [^8] Our dataset is narrower, considering only loans accompanied by textual descriptions, as detailed in Section 4., which aligns with expectations in the credit risk context.
@@
Figure 7 reveals a direct and linear relationship between the BERT score and the SHAP values, which in this case relate to the default risk. Notably, this relationship is asymmetric around BERT score values of 0.5; scores below 0.4 correspond to SHAP values ranging between -0.4 and -0.6, strongly guiding the model toward predicting non-default. Conversely, this impact range in the positive case is only reached by BERT scores exceeding 0.7, signifying that only exceptionally high BERT scores serve as strong indicators of default.
7.3 A closer examination of the role of the BERT score in classification
(a) Loan (correctly) classified as non-default.
(b) Loan (correctly) classified as default.
Figure 8: Waterfall plots representing the SHAP values of the features in two loans.
As shown in our analysis of feature importance in XGBoost, the BERT score plays a crucial role in the correct classification of the loans. Using SHAP values, we can determine how each variable influences the classification decision for a given loan. Figure 8 shows the waterfall plot for two loans where the BERT score is a decisive factor. The plot should be read from bottom to top. It begins at the bottom with the expected value of the model output. Each row then shows how each feature’s positive (red) or negative (blue) contribution shifts the prediction from this expected value to the final prediction for that specific loan. Features are ordered from bottom to top in ascending order of contribution (in absolute value). The x-axis represents log-odds units (the margin output before the logistic link function used by the classifier), meaning that negative values correspond to probabilities below 0.5 for classifying the loan as default.
Figure 8(a) shows a loan with a low BERT score (0.37) and a SHAP value of $-0.47$, the highest in absolute value for that instance. This SHAP value counteracts the influence of other variables, leading to a (correct) classification of the loan as non-default. Conversely, Figure 8(b) illustrates the opposite scenario: a loan with a high BERT score (0.686), where the SHAP value—again, the highest for that instance—overrides the effect of other variables and correctly assigns the loan to the default class. Below are the description of both loans given by its borrowers:
Purpose: Home improvement; BERT score: 0.37
“I want to have an in ground pool built with an aluminum screened enclosure. Nothing super fancy, just enough to enjoy with my wife a kids.”
Purpose: Debt consolidation; BERT score: 0.686
“I have problems with one of my loans what I have it is with household-beneficial finance bank it is for \$15,250.00 and the interest rate is 25.8% and the payment is 323 monthly and I have 3 months pass due, I page all time in time but the situacion make loss that way cause the gas bill was very higher and the last mont the second day my Mother die and was very expensive for us her funeral cause she die here but we bring her body to Nicaragua, and the rest is for a loan with American General Finance \$2,500.00, Wells Fargo \$480.00 , Fifth Third Bank Optimun \$357.00, all make a total of \$18,587.00, please I need cause I get some problems with the big loan with household-beneficial.”
Although Section 6.2 suggested that shorter texts tend to have higher BERT scores, this is not the case here. BERT’s ability to analyze content and textual characteristics enables it to correctly associate the long text with a high probability of default and the shorter one with a higher likelihood of repayment.
@@ 7.3 A closer examination of the role of the BERT score in classification
7.4 Impact of the BERT score in the classification results across the purpose categories
@@
In this section, we examine how incorporating the BERT score in the model changed classification outcomes across various categories of the purpose variable. Given the language comprehension capabilities of LLMs, it is conceivable that the BERT score provides a more nuanced characterization of the risk associated to the loan purposes than the categorical purpose variable alone. For instance, the inherently ambiguous ‘other’ category may benefit from the nuanced understanding of loan descriptions provided by the BERT model, potentially leading to improved prediction outcomes.
@@
Figure 9: BACC changes (in relative percentage) by purpose after including the BERT score.
@@
Figure 9 illustrates the relative changes in balanced accuracy for each loan purpose after adding the BERT score. Notably, while the ‘other’ category shows a modest improvement of 2.11%, several other categories exhibit more significant enhancements, including ‘educational’ (9.22%), ‘moving’ (5.84%), ‘medical’ (3.84%), and ‘small business’ (3.40%). Given the black-box nature of our model, it is not easy to ascertain why these categories have improved more than the others. However, we posit that these categories share a commonality—the more detailed specification of purpose or a deeper understanding of the borrower’s situation contributes to a more precise delineation of the default risk. For instance, ‘education’ loans might incorporate information about the borrower’s field of study or educational institution, which could correlate with employability and repayment capacity. Similarly, in categories like ‘moving’, ‘medical’, or ‘small business’ loans, the BERT score likely reflects a deeper understanding of the borrower’s situation, enabling a more accurate default risk assessment.
@@
The following examples illustrate instances of loans that were initially misclassified as defaults without the BERT score but were correctly predicted as non-defaults after its inclusion:
@@
In these cases, the corresponding BERT scores are consistently below 0.4. As shown in Figure 7, scores below this threshold are strong drivers for predicting non-default. All of these texts offer precise descriptions of the purpose of the loan or the borrower’s situation, which allows a relatively moderate risk to be anticipated.
(a) Distributions of the BERT score.
(b) Distributions of the SHAP values.
Figure 10: BERT score and SHAP values distributions for the educational (orange), moving (green), and the rest of purposes (blue).
To better understand the differences in classifier performance across purposes, we conduct an additional analysis focusing on the purposes with the greatest performance improvement. Figure 10(a) shows the distributions of BERT score values for loans with educational and moving purposes, compared to the distribution of loans with all other purposes, which serves as the reference. We can see that the reference distribution is more sharply peaked, indicating that most loans have average BERT score values that contribute little information to the classification. In contrast, the distribution of the loans with an educational purpose is shifted to the left and exhibits a higher density of low BERT score values (below 0.4), which indicates that the loan will likely be repaid. In the case of the moving purpose, the most pronounced difference can be seen in the right tail, as it has a higher frequency of values above 0.6, which helps to detect loans likely to default.
Table 14: Coefficients of the quantile regression at the 5th, 10th, 50th, 90th, and 95th percentiles.
5th 10th 50th 90th 95th
Intercept 0.4101 0.4328 0.5178 0.5890 0.6106
Default +0.0164* +0.0166* +0.0127* +0.0124* +0.0146*
Educational --0.0208* --0.0240* --0.0205* --0.0308* --0.0316*
Moving --0.0022 --0.0014 +0.0002 --0.0027 +2.8e-05
Default–Educational --0.0067 --0.0077 +0.0126 +0.0288 +0.0084
Default–Moving --0.0085 --0.0040 +0.0062 +0.0286* +0.0357*
* p-value less than 0.01.
To assess the significance of these differences, we perform a quantile regression with interactions at five percentiles: the median (50th) and the two tails of the distribution (5th, 10th, 90th and 95th). The quantile regression model is specified as:
[…]
Here, $D_{E}$ and $D_{M}$ represent the dummy variables for educational and moving loans, respectively, with all other purposes serving as the reference group. Meanwhile, $\text{Default}_{i}$ is an indicator variable that denotes whether the loan eventually defaulted or not. The quantile regression allows us to assess whether there are significant differences in the BERT score between defaulted and non-defaulted moving and educational loans compared to loans of other purposes.
The results are presented in Table 14. At the 5th, 10th, 50th, 90th and 95th percentiles, the intercept values (0.4101, 0.4328, 0.5178, 0.5890, and 0.6106) capture the baseline BERT score for loans that do not belong to the educational or moving purposes and that were returned (i.e., $D_{E}=0$, $D_{M}=0$, and $\text{Default}_{i}=0$). The positive and statistically significant coefficients for default across all quantiles indicate that default is associated with an increase in BERT score (+0.0164, +0.0166, +0.0127, +0.0124, and +0.0146). Thus, the first conclusion that we can draw is that our fine-tuned BERT successfully assigns a higher risk to the loans that eventually defaulted, at least in the quantiles analyzed, in line with the results obtained in previous analyses.
The educational dummy variable shows significant negative coefficients across all quantiles (--0.0208, --0.0240, --0.0205, --0.0308, and --0.0316), confirming that, regardless of default status, educational loans have lower BERT scores at the quantiles analyzed compared to the reference group—consistent with a lower risk profile observed in Figure 10(a). Regarding the interaction term, the Default–Educational coefficient is not significant at any quantile, suggesting that defaulted educational loans do not have significantly different BERT scores compared to repaid educational loans.
In contrast, the coefficients of the moving dummy when the default variable is zero are not statistically significant at any quantile. This implies that, in such cases, non-defaulted moving loans do not differ from the non-defaulted reference group in their baseline BERT scores. As for variable interaction, the Default–Moving coefficient is significant at the 90th and 95th percentiles (+0.0286 and +0.0357). This result indicates that, at the upper end of the BERT score distribution, defaulted loans from the moving purpose have higher BERT scores than defaulted loans from the rest of the purposes. In other words, the difference observed in Figure 10(a) in the right tail of the BERT score distribution for moving loans relative to the reference group is driven by the higher risk associated with defaulted loans.
Overall, these findings evidence a nuanced relationship: default increases BERT scores across the board, non-defaulted educational loans exhibit lower scores—reflecting lower risk—than other categories, and moving loans that eventually defaulted exhibit a higher risk reflected only in the upper quantiles. To conclude the analysis, Figure 10(b) shows the distribution of SHAP values for educational and moving purposes compared to the rest of the purposes. Again, educational loans show a distribution shifted to the left and with a denser left tail, indicating that XGBoost uses this variable to identify loans likely to be repaid. In the case of moving loans, differences are observed at the extremes, especially in the right tail, indicating the improvement in discrimination that this variable provides to the classifier. This analysis shows that the interaction between the BERT score and the educational and moving purposes is the reason behind the remarkable performance improvements observed in these categories.
As a conclusion, while previous research had reported mixed results regarding the predictive power of linguistic factors for loan default [16, 54, 15, 46], our findings suggest that an LLM-based risk indicator used at the granting stage has a significant positive influence in the classification results, thus demonstrating BERT’s capacity to draw meaningful information from the loan descriptions.
8 Conclusion
In this paper, we presented a novel approach that leverages state-of-the-art natural language processing techniques to enhance credit risk models. By fine-tuning BERT on loan descriptions, we generated a risk score that effectively distinguishes between defaulted and non-defaulted loans, particularly at the granting stage, when decision-makers have limited variables available to inform their decisions. In addition, integrating this BERT-based risk score with traditional variables significantly improved the performance of conventional loan-granting models. This result aligns with those obtained by Xia et al. [57]. Our analysis suggests that the information extracted by the language model can capture aspects of the text related to the linguistic aspects but also with content-based factors related to loan purpose and the borrower’s creditworthiness. The inclusion of the BERT-based risk score also reshapes the classifier’s decision-making process and the role played by other variables.
Our approach can be easily applied without the need for manual annotation, which is a complex and subjective task. Additionally, while fine-tuning the LLM is a computationally intensive process that requires GPU resources, generating predictions such as the risk score for a loan description can be done rapidly on standard home equipment, making this approach highly accessible. This work opens several avenues for further exploration to refine both the predictive capability of the model and our understanding of loan applicants’ situations.
In our work, we have thoroughly documented the model’s architecture as well as the preprocessing, fine-tuning, and training procedures. We have also used SHAP values to understand how the final model classifies and even how it makes decisions for individual instances. While these aspects are crucial in financial applications, a key limitation remains: the lack of transparency in how the BERT-based risk score is generated, which limits the understanding of the factors influencing the score and its potential biases. This challenge hinders the practical application of this approach in real-world settings, where it is critical to understand the score generation process and ensure that it does not introduce biases related to gender, ethnicity, or financial inclusion. Therefore, enhancing the transparency and explainability of the BERT score is essential, not only for regulatory compliance but also to build trust among borrowers and lenders [43].
In this respect, several approaches exist to better understand how neural-based Natural Language Processing models such as BERT work [31]. Intrinsic methods, such as inspecting attention weights, do not offer sufficient transparency or meaningful explanations for model predictions [21, 45]. Still, it is possible to use surrogate models, such as LIME [41] or anchor rules [42], to try to interpret BERT-based predictions, as has been done in fake news detection [51]. Such an approach can help to better understand how these models work and facilitate their adoption in real-world settings.
@@ 8 Conclusion
A different approach that could also be explored is the use of LLM-based topic modeling techniques, such as BERTopic [18]. It could help to identify “risky topics” by making use of the deep semantic understanding of LLMs. These topics could then be used as additional input variables in the granting model, improving both predictive performance and transparency in the decision-making process. Additionally, the embeddings from BERT or other LLMs could be used to generate interpretable topics from large, complex vocabularies, as demonstrated in other studies [14]. Applying such topic modeling to loan descriptions could help identify loans with varying risk levels, further enhancing the transparency of credit risk assessments.
@@
Further research could also explore the use of advanced LLMs. Encoder-only models like RoBERTa [28] may capture more intricate linguistic patterns while emerging generative AI approaches could redefine how risk scoring is performed. Although these models were not originally designed for such tasks, recent advancements have shown promising approximations. For instance, techniques like CARP (Clue And Reasoning Prompting) [49] utilize in-context learning with few-shot examples to perform classification without fine-tuning. Applying such methods to risk scoring may open new possibilities for achieving robust results while minimizing computational overhead.
@@
Future work should also address the economic implications of our findings by integrating cost- or profit-sensitive approaches [56, 3]. Investigating how the inclusion of textual descriptions impacts financial outcomes could provide valuable insights into the practical utility of these methods. By linking improved prediction performance to tangible economic benefits, we can further bridge the gap between academic innovation and real-world application.
Acknowledgements
We thank Antonio Caparrini López and Miller Janny Ariza Garzón for their help at different stages of the project.